2 π r^2 + 2 π r h
Solve for h.

Answer by Boreal(15235)

Need an equation. This is the area of a closed cylinder.A=2 pi*r^2+2 pi *rh
A=2pir(r+h)
A/2pi*r=(r+h)
h=(A/2pir)-r
There are other ways to write this but this is reasonable.